# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ Abstract
We study weighted pseudorandom generators (WPRGs) and derandomizations for read-once branching programs (ROBPs), which are key problems towards answering the fundamental open question $\mathbf{BPL}\stackrel{{\scriptstyle?}}{{=}}\mathbf{L}$. Denote $n$ and $w$ as the length and the width of a ROBP. We have the following results.
For standard ROBPs, there exists an explicit $\varepsilon$-WPRG with seed length
$O\left(\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w\left(\log\log\log w-\log\log\max\left\{2,\frac{\log w}{\log n/\varepsilon}\right\}\right)+\log(1/\varepsilon)\right).$
When $n=w^{o(1)}$, this is better than the constructions in [Hoz21, CDR+21, PV21, CL20]. Further, by using this result in a black-box way, we attain a WPRG for regular ROBPs with seed length
This slightly improves the result of [CHL+23] and matches the recent simultaneous and independent work of [CTS25].
For permutation ROBPs with unbounded widths and single accept nodes, there exists an explicit $\varepsilon$-WPRG with seed length
$O\left(\log n\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)+\log(1/\varepsilon)\right).$
This slightly improves the result of [CHL+23].
For regular ROBPs with width $w$, length $n\leq 2^{O(\sqrt{\log w})},\varepsilon=1/\operatorname{\mathrm{poly}}w$, and multiple accept nodes, we give the first derandomization which attains optimal space complexity $O(\log w)$, i.e. in $\mathbf{L}$ exactly. When requiring the derandomization in $\mathbf{L}$, the only previous method that can work is using the Nisan-Zuckerman generator [NZ96] or its variations in [Arm98]. Our result significantly improves the length and error of the regular ROBPs that can be derandomized under this requirement. Also, if compared to the derandomization of [AKM+20, CHL+23, CL24], then our result not only improves the space complexity to optimal, but also improves the time complexity from super-polynomial to standard polynomial.
All our results are based on iterative applications of weighted pseudorandom reductions, which is a natural new notion for simplifying ROBPs.
@@ 1 Introduction
Using the probabilistic method, one can show the existence of a non-uniform $\varepsilon$-PRG for standard-order ROBPs of length $n$, width $w$, and alphabet size $2^{s}$, with an optimal seed length of $O(s+\log(nw/\varepsilon))$. However, constructing explicit PRGs that have short seed lengths turns out to be an exceptional challenge. In a seminal work, Nisan [Nis92a] constructed an explicit $\varepsilon$-PRG for ROBPs of length $n$, width $w$, and binary alphabet $\{0,1\}$, with a seed length of $O(\log n\log(nw/\varepsilon))$. Building on this result, Saks and Zhou [SZ99] developed a celebrated algorithm to derandomize $\mathbf{BPL}$, within $O(\log^{3/2}n)$ space deterministically. Nisan [Nis92b] also used [Nis92a] to show that $\mathbf{BPL\subseteq\mathbf{SC}}$. Impagliazzo, Nisan, and Wigderson [INW94] generalized the construction of [Nis92a] by using expanders, to fool more general models in network communication. In the meantime, for short ROBPs, Nisan and Zuckerman [NZ96] gave another remarkable PRG that has a better seed length for wide but short ROBPs, i.e. with seed length $O(\log w)$ when $n=\operatorname{\mathrm{poly}}\log w,\varepsilon=2^{-\log^{0.99}w}$. Armoni [Arm98] further extended [Nis92a, NZ96] to construct an improved PRG with seed length333[Arm98] needs to use an extractor with seed length optimal up to constant factors which is discovered later than [Arm98], e.g. [GUV09]. $O\left(\frac{\log n\log(nw/\varepsilon)}{\max\{1,\log\log w-\log\log(n/\varepsilon)\}}\right)$.
@@ 1.1 Weighted Pseudorandom Generators
Despite years of research, the challenging problem of constructing better PRGs for general ROBPs remains open. Nisan’s generator is still the best explicit PRG for ROBPs when lengths and widths are large, e.g. $n=w=1/\varepsilon$. However, PRGs are not the only black-box solution to derandomization. Weighted pseudorandom generator (WPRG) and Hitting set generator (HSG) can also do derandomizations.
A remarkable work by Braverman, Cohen, and Garg [BCG19] improves the seed length by introducing and constructing WPRGs.
@@
Let $\mathcal{F}$ be a class of ROBPs $f:(\{0,1\}^{s})^{n}\to\{0,1\}$. A $W$-bounded $\varepsilon$-WPRG for $\mathcal{F}$ is a function $(G,w):\{0,1\}^{d}\to(\{0,1\}^{s})^{n}\times\mathbb{R}$ such that for every $f\in\mathcal{F}$, we have
$\left|\mathbb{E}_{x\in(\{0,1\}^{s})^{n}}f(x)-\sum_{r\in\{0,1\}^{d}}\left[\frac{1}{2^{d}}w(r)\cdot f(G(r))\right]\right|\leq\varepsilon,$
$\forall r,|w(r)|\leq W.$
@@
As shown in [BCG19], under this notion, the seed length can be significantly improved to $\widetilde{O}(\log n(\log n+\log w)+\log\frac{1}{\varepsilon})$, i.e. for $\varepsilon$ there is only an isolated addend. Chattopadhyay and Liao [CL20] further improved this isolated addend to be the optimal $O(\log(1/\varepsilon))$. Then Cohen, Doron, Renard, Renard, Sberlo, and Ta-Shma [CDR+21], and also Pyne and Vadhan [PV21] attained black-box error reductions from PRGs to WPRGs, which can also achieve an isolated error term in the seed length. They both are based on a preconditioned Richardson Iteration, which was previously developed by Ahmadinejad, Kelner, Murtagh, Peebles, Sidford, and Salil Vadhan [AKM+20] for high precision derandomization of random walks on Eulerian graphs. Hoza [Hoz21] further improved the WPRG construction to attain seed length $O(\log n\log(nw)+\log(1/\varepsilon))$.
@@
HSG is an even weaker notion than WPRG. However, it is also a powerful tool in derandomization.
@@
One can use HSG to find an accepting path when the acceptance probability is significant. Actually, HSG is more powerful than this. Cheng and Hoza [CH22] showed how to approximate the acceptance probabilities of ROBPs of length $n$, width $w$ using an HSG for ROBPs with significantly larger width $O(\operatorname{\mathrm{poly}}(nw/\varepsilon))$ and length $O(\operatorname{\mathrm{poly}}(nw/\varepsilon))$. Pyne, Raz, and Zhan [PRZ23] further extended the method to give a deterministic sampler for such a task. Constructing better HSGs is also a challenging task. Hoza and Zuckerman gave the current best HSG with seed length $O\left(\frac{\log n\log(nw)}{\max\{1,\log\log w-\log\log n\}}+\log(1/\varepsilon)\right)$ for binary alphabet.
@@ 1.2 WPRG for short-wide standard ROBPs
Although significant progress has been made in the study of WPRGs for general ROBPs with large length and width, for shorter lengths, the existing work still do not match the seed length of the HSG construction [HZ20]. This naturally raises the question: can we construct a WPRG for shorter but wide ROBPs with better seed length? Note that for the special case $n=\operatorname{\mathrm{poly}}\log w$, NZ PRG [NZ96] and Armoni’s PRG [Arm98] have optimal seed lengths of $O(\log w)$ when $\varepsilon=2^{-\log^{1-v}w}$, where $v$ is any constant in $(0,1)$. However when the error is smaller such as $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$, the seed lengths of the WPRGs of [CDR+21, Hoz21, PV21] deteriorate to $O(\log w\log\log w)$, which have no advantage than those of the PRGs [Nis92a, INW94, Arm98].
@@
$O\left(\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w\left(\log\log\log w-\log\log\max\left\{2,\frac{\log w}{\log n/\varepsilon}\right\}\right)+\log(1/\varepsilon)\right)$
@@
This WPRG has a better seed length than Armoni’s PRG [Arm98, KNW08]. Also, this slightly outperforms [Hoz21, CDR+21, PV21, CL20], when $n\ll w$.
For the Nisan-Zuckerman regime i.e. $n=\operatorname{\mathrm{poly}}\log w$, $\varepsilon=2^{-\log^{1-\alpha}n}$ for any constant $\alpha>0$, one can see that the seed length of our Theorem 1.5 is $O(\log w)$, by noting that the subtraction of the two triple-log terms becomes a constant. This matches the seed length of the Nisan-Zuckerman PRG [NZ96]. Further when $n$ is larger or $\varepsilon$ is smaller, our seed length is strictly better.
Table 1 and Table 2 summarize the seed length of $\varepsilon$-PRGs and $\varepsilon$-WPRGs for short-wide general ROBPs.
@@
$O\left(\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log(w/\varepsilon)\log\log_{n}(1/\varepsilon)\right)$ WPRG [CDR+21]
@@
Table 1: Comparison of seed length of $\varepsilon$-PRGs, $\varepsilon$-HSGs and $\varepsilon$-WPRGs for general ROBPs of width $w$, length $n<w$, and alphabet $\{0,1\}$, where $\varepsilon\leq 1/\operatorname{\mathrm{poly}}(w)$.
@@
$O(\log(w/\varepsilon)\log\log w)$ PRG [Arm98],[INW94],[Nis92a] $\varepsilon\leq 1/\operatorname{\mathrm{poly}}(w)$
$O(\log w\log\log w+\log(1/\varepsilon))$ WPRG [Hoz21],[CDR+21] $\varepsilon\leq 1/\operatorname{\mathrm{poly}}(w)$
@@
Table 2: Comparison of seed length of $\varepsilon$-PRGs, $\varepsilon$-HSGs and $\varepsilon$-WPRGs for short-wide ROBPs of width $w$, length $n=\operatorname{\mathrm{poly}}(\log w)$, and alphabet $\{0,1\}$.
One direct application of our new WPRG is to improve the WPRG for regular ROBPs.
@@
Chen, Hoza, Lyu, Tal, and Wu [CHL+23] gave a WPRG for regular ROBPs. Their seed length is $\widetilde{O}\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)+\log(1/\varepsilon)\right).$ Using our WPRG instead of the INW PRG in one step of their construction, we can eliminate the low-order multiplicative factors in their seed length, which are double-logarithmic multiplicative factors attached to both $\log(1/\varepsilon)$ and $\log n\sqrt{\log(1/\varepsilon)}$. Our result is stated as the following.
@@
For every $w,n\in\mathbb{N}$ with $n\geq\log\log w$, and every $\varepsilon\geq 2^{-n}$, there exists an explicit $\varepsilon$-WPRG with seed length
$O\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)+\log(1/\varepsilon)\right)$
for regular ROBPs of width $w$, length $n$ and alphabet $\{0,1\}$.
A recent simultaneous and independent work of Chen and Ta-Shma [CTS25] also achieves this seed length. Their construction combines the work of [CHL+23] and Hoza [Hoz21]. On the contrary, our construction is using our Theorem 1.5 to replace a step of using INW PRG in [CHL+23].
1.3 WPRG for unbounded-width permutation ROBPs with a single accepting node
Permutation ROBPs are interesting special ROBPs, where the transition functions between layers are permutations.
@@ 1.3 WPRG for unbounded-width permutation ROBPs with a single accepting node
Early work on PRGs for permutation ROBPs [BV10, De11, KNP11, Ste12, RSV13, CHHL19] focus on the constant width case. Inspired by the progress on derandomizing squares [RV05, AKM+20, APP+23], recent studies focus on unbounded-width permutation ROBPs with a single accepting node [HPV21a, PV21, BHPP22, CHL+23]. Hoza, Pyne and Vadhan [HPV21a] designed an $\varepsilon$-PRG for unbounded-width permutation ROBPs with seed length $\widetilde{O}(\log n\log(1/\varepsilon))$ for binary alphabet. They also proved that any PRG for this class must have seed length $\widetilde{\Omega}(\log n\log(1/\varepsilon))$. For the WPRG case, Pyne and Vadhan [PV21] designed a $\varepsilon$-WPRG with seed length $\widetilde{O}(\log n\sqrt{\log(n/\varepsilon)}+\log(1/\varepsilon))$. This was improved by Chen, Hoza, Lyu, Tal, and Wu [CHL+23] to [^4] The detailed parameter of [CHL+23] is $O\left(\log n\sqrt{\log(1/\varepsilon)}\sqrt{\log\log(n/\varepsilon)}+\log(1/\varepsilon)\log\log(n/\varepsilon)\right)$. $\widetilde{O}(\log n\sqrt{\log(1/\varepsilon)}+\log(1/\varepsilon))$.
Our idea in WPRG construction for general ROBPs can also be used to slightly improve the WPRG construction for permutation ROBPs. Specifically, we have the following result.
@@
for the class of permutation ROBPs of length $n$ and alphabet $\{0,1\}^{s}$ with a single accepting node.
When the error is small enough, i.e. $\varepsilon\leq 2^{-\log^{2}n}$, our WPRG has a seed length optimal up to a constant factor. The comparison of our result with the previous results is shown in Table 3.
@@
1.4 Derandomization for short regular ROBPs
There is an extensive body of work focused on constructing HSGs, WPRGs, and PRGs for regular ROBPs [BRRY14, BHPP22, De11, RSV13, CL24, CHL+23]. These results also have shed light on derandomizing standard ROBPs and space-bounded computations.
PRGs and WPRGs constructed for regular ROBPs naturally induce derandomization algorithms for this class. However, dedicated derandomization algorithms achieve better performance. Ahmadinejad, Kelner, Murtagh, Peebles, Sidford, and Vadhan [AKM+20] derandomized regular ROBPs of length $n$, width $w$ and alphabet $\{0,1\}$ within error $\varepsilon$, using space $\widetilde{O}(\log(nw)\log\log(1/\varepsilon))$. Subsequently, Chen, Hoza, Lyu, Tal, and Wu [CHL+23] achieved the same result with a simplified algorithm. Chattopadhyay and Liao [CL24] constructed another alternate algorithm with the same space complexity.
Our next result focus on derandomization algorithms for short-wide regular ROBPs and constructs a derandomization algorithm with improved space complexity and time complexity for this regime, by adapting our WPRG for permutation ROBPs.

### v3

@@ Abstract
We study weighted pseudorandom generators (WPRGs) and derandomizations for read-once branching programs (ROBPs). Denote $n$ and $w$ as the length and the width of a ROBP. We have the following results.
For standard ROBPs, we give an explicit $\varepsilon$-WPRG with seed length
$O\left(\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w\left(\log\log\log w-\log\log\max\left\{2,\frac{\log w}{\log\frac{n}{\varepsilon}}\right\}\right)+\log\frac{1}{\varepsilon}\right).$
When $n=w^{o(1)}$, this is better than the WPRGs of [Hoz21, CDR+21, PV21, CL20]. Further as a direct application, we attain a WPRG for regular ROBPs with a better seed length than that of [CHL+23, CL24].
For permutation ROBPs with unbounded widths and single accept nodes, we give an explicit $\varepsilon$-WPRG with seed length
$O\left(\log n\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)+\log(1/\varepsilon)\right),$
improving [CHL+23]. A key difference to [CHL+23] is that this implies a WPRG with optimal seed length for short-wide ROBPs with multiple accept nodes. Specifically, after switching to multiple accept nodes in a standard way by replacing $\varepsilon$ with $\varepsilon/w$, this gives a WPRG with optimal seed length $O(\log w)$ for $n=2^{O(\sqrt{\log w})}$, and error $1/\operatorname{\mathrm{poly}}w$. The only previous work attaining optimal seed lengths are Nisan-Zuckerman style PRGs [NZ96, Arm98] but they are only optimal for $n=\operatorname{\mathrm{poly}}\log w,\varepsilon=2^{-\log^{0.9}w}$.
We also give a new Nisan-Zuckerman style derandomization for regular ROBPs with width $w$, length $n=2^{O(\sqrt{\log w})}$, and multiple accept nodes. We attain optimal space complexity $O(\log w)$ for arbitrary approximation error $\varepsilon=1/\operatorname{\mathrm{poly}}w$. When requiring the derandomization to be in $\mathbf{L}$, again the only previous result is by Nisan-Zuckerman style PRGs [NZ96, Arm98], which are only optimal for $n=\operatorname{\mathrm{poly}}\log w,\varepsilon=2^{-\log^{0.9}w}$. Also, if compared to [AKM+20, CHL+23, CL24], which can be viewed as Saks-Zhou style derandomizations, then for $n=2^{O(\sqrt{\log w})}$ our derandomization not only improves the space complexity to optimal, but also substantially improves the time complexity from super polynomial to standard polynomial in $w$. Note that derandomizations of [AKM+20, CHL+23, CL24] has space complexity $S=O(\log(nw)\log\log(nw/\varepsilon))$ and time complexity exponential in $S$.
All our results are based on iterative weighted pseudorandom reductions, which can iteratively reduce fooling long ROBPs to fooling short ones.
@@ 1 Introduction
By the probabilistic method, one can show the existence of a non-uniform $\varepsilon$-PRG for standard-order ROBPs of length $n$, width $w$, and alphabet size $2^{s}$, with an optimal seed length of $O(s+\log(nw/\varepsilon))$. However, constructing explicit PRGs that have short seed lengths turns out to be an exceptional challenge. In a seminal work, Nisan [Nis92a] constructed an explicit $\varepsilon$-PRG for ROBPs of length $n$, width $w$, and binary alphabet $\{0,1\}$, with a seed length of $O(\log n\log(nw/\varepsilon))$. Building on this, Saks and Zhou [SZ99] developed a celebrated algorithm to derandomize $\mathbf{BPL}$, within $O(\log^{3/2}n)$ space deterministically. Nisan [Nis92b] also used [Nis92a] to show that $\mathbf{BPL\subseteq\mathbf{SC}}$. Impagliazzo, Nisan, and Wigderson [INW94] generalized the construction of [Nis92a] by using expanders, to fool more general models in network communication. In the meantime, for short ROBPs, Nisan and Zuckerman [NZ96] gave another remarkable PRG that has an optimal seed length for wide but short ROBPs, i.e. with seed length $O(\log w)$ when $n=\operatorname{\mathrm{poly}}\log w,\varepsilon=2^{-\log^{0.99}w}$. Armoni [Arm98] further extended [Nis92a, NZ96] to construct an improved PRG with seed length333[Arm98] needs to use an extractor with seed length optimal up to constant factors which is discovered later than [Arm98], e.g. [GUV09]. $O\left(\frac{\log n\log(nw/\varepsilon)}{\max\{1,\log\log w-\log\log(n/\varepsilon)\}}\right)$.
@@ 1.1 Weighted Pseudorandom Generators
Despite years of research, the challenging problem of constructing better PRGs for general ROBPs remains open. However, PRGs are not the only black-box method for derandomizations. A remarkable work by Braverman, Cohen, and Garg [BCG19] improves the seed length by introducing and constructing WPRGs.
@@
Let $\mathcal{F}$ be a class of ROBPs $f:(\{0,1\}^{s})^{n}\to\{0,1\}$. A $W$-bounded $\varepsilon$-WPRG for $\mathcal{F}$ is a function $(G,\sigma):\{0,1\}^{d}\to(\{0,1\}^{s})^{n}\times\mathbb{R}$ such that for every $f\in\mathcal{F}$, we have
$\left|\mathbb{E}_{x\in(\{0,1\}^{s})^{n}}f(x)-\sum_{r\in\{0,1\}^{d}}\left[\frac{1}{2^{d}}\sigma(r)\cdot f(G(r))\right]\right|\leq\varepsilon,$
$\forall r,|\sigma(r)|\leq W.$
@@
As shown in [BCG19], under this notion, the seed length can be significantly improved to $\widetilde{O}(\log n\log(nw)+\log\frac{1}{\varepsilon})$, i.e. for $\varepsilon$ there is only an isolated addend. Chattopadhyay and Liao [CL20] further improved this construction to attain seed length $O(\log n\log(nw)\log\log(nw)+\log\frac{1}{\varepsilon})$. Then Cohen, Doron, Renard, Renard, Sberlo, and Ta-Shma [CDR+21], and also Pyne and Vadhan [PV21] gave black-box error reductions from large error PRGs to small error WPRGs. They both are based on a preconditioned Richardson Iteration, which was previously developed by Ahmadinejad, Kelner, Murtagh, Peebles, Sidford, and Salil Vadhan [AKM+20] for high precision derandomization of random walks on Eulerian graphs. Hoza [Hoz21] further improved the error reduction based WPRG constructions to attain seed length $O(\log n\log(nw)+\log(1/\varepsilon))$.
@@
WPRG is stronger than Hitting Set Generator (HSG), while HSG is already a powerful tool in derandomization.
@@
One can use HSG to find an accepting path when the acceptance probability is significant. Actually, HSG is more powerful than this. Cheng and Hoza [CH22] showed how to approximate the acceptance probabilities of ROBPs by HSGs for larger-size ROBPs. Pyne, Raz, and Zhan [PRZ23] further extended the method to give a deterministic sampler for such tasks. Constructing better HSGs is also a challenging task. Hoza and Zuckerman gave a HSG with seed length $O\left(\frac{\log n\log(nw)}{\max\{1,\log\log w-\log\log n\}}+\log(1/\varepsilon)\right)$ which is optimal when $n=\operatorname{\mathrm{poly}}\log w$.
@@ 1.2 WPRG for short-wide standard ROBPs
Nisan-Zuckerman PRG and Armoni’s generator are better in seed length than Nisan’s PRG or INW PRG for short-wide cases when the error is large. It is a natrual question whether one can transform these PRGs to WPRGs to attain better seed lengths for smaller errors. A specific interesting case is $n=\operatorname{\mathrm{poly}}\log w$, for which NZ PRG [NZ96] and Armoni’s PRG [Arm98] have optimal seed lengths of $O(\log w)$ when $\varepsilon=2^{-\log^{1-v}w}$, where $v$ is any constant in $(0,1)$. However when the error is smaller such as $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$, the seed lengths of the WPRGs of [CDR+21, Hoz21, PV21] deteriorate to $O(\log w\log\log w)$, which have no advantage than those of the PRGs [Nis92a, INW94, Arm98].
@@
$O\left(\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}+\log w\left(\log\log\log w-\log\log\max\left\{2,\frac{\log w}{\log(n/\varepsilon)}\right\}\right)+\log(1/\varepsilon)\right)$
@@
The seed length is better than [Hoz21, CDR+21, PV21, CL20], when $n\ll w$. For the Nisan-Zuckerman regime i.e. $n=\operatorname{\mathrm{poly}}\log w$, $\varepsilon=2^{-\log^{1-\alpha}n}$ for any constant $\alpha>0$, the seed length of our Theorem 1.5 is $O(\log w)$, since the subtraction of the two triple-log terms becomes a constant. This matches the seed length of the Nisan-Zuckerman PRG [NZ96]. Further when $n$ is larger or $\varepsilon$ is smaller, our seed length is strictly better. Table 1 and Table 2 summarize the seed length of $\varepsilon$-PRGs and $\varepsilon$-WPRGs for short-wide general ROBPs.
@@
$\widetilde{O}(\log n\log(nw)+\log\frac{1}{\varepsilon})$ WPRG [BCG19]
$O(\log n\log(nw)\log\log(nw)+\log\frac{1}{\varepsilon})$ WPRG [CL20]
$O\left(\log n\log(nw)+\log(w/\varepsilon)\log\log_{n}(1/\varepsilon)\right)$ WPRG [CDR+21]
@@
Table 1: Comparison of seed length of $\varepsilon$-WPRGs for general ROBPs of width $w$, length $n<w$, and alphabet $\{0,1\}$.
@@
$O(\log(w/\varepsilon)\log\log w)$ PRG [Arm98, INW94, Nis92a] $\varepsilon\leq 1/\operatorname{\mathrm{poly}}(w)$
$O(\log w\log\log w+\log(1/\varepsilon))$ WPRG [Hoz21, CDR+21] $\varepsilon\leq 1/\operatorname{\mathrm{poly}}(w)$
@@
Table 2: Comparison of seed length of $\varepsilon$-WPRGs for short-wide ROBPs of width $w$, length $n=\operatorname{\mathrm{poly}}(\log w)$, and alphabet $\{0,1\}$.
Another interesting point is that our construction deploys a new framework which iteratively applies weighted pseudorandom reductions. We will describe this in detail in later sections.
As a direct application of our WPRG, we also attain a better WPRG for regular ROBPs.
@@
For every $w,n\in\mathbb{N}$, and every $\varepsilon>0$, there exists an explicit $\varepsilon$-WPRG with seed length
for regular ROBPs of width $w$, length $n$ and binary alphabet.
This is better than that of Chen, Hoza, Lyu, Tal, and Wu [CHL+23] which has a seed length $\widetilde{O}\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)+\log(1/\varepsilon)\right)$, and matches an independent work of Chen and Ta-Shma [CTS25] for the binary case. For larger alphabets, we can actually attain a seed length with a better dependence on the alphabet size than [CTS25]. See our Theorem 3.19 for details. The construction of [CTS25] combines the methods of [CHL+23] and Hoza [Hoz21]. Our construction is using our main technical theorem Theorem 3.2 to replace the INW PRG generating correlated seeds in [CHL+23], with an argument based on weighted pseudorandom reductions.
1.3 WPRG for unbounded-width permutation ROBPs with a single accept node
Our framework works even better for constructing WPRGs against permutation ROBPs. Permutation ROBPs are interesting special ROBPs, where the transition functions between layers are permutations.
@@ 1.3 WPRG for unbounded-width permutation ROBPs with a single accept node
Early work on PRGs for permutation ROBPs [BV10, De11, KNP11, Ste12, RSV13, CHHL19] focus on constant width cases. A remarkable line of recent studies develops PRG/WPRG/HSGs for unbounded-width permutation ROBPs with single accepting nodes [HPV21a, PV21, BHPP22, CHL+23]. Hoza, Pyne and Vadhan [HPV21a] showed an $\varepsilon$-PRG for unbounded-width permutation ROBPs with seed length $\widetilde{O}(\log n\log(1/\varepsilon))$. They also proved that any PRG for this class must have seed length $\widetilde{\Omega}(\log n\log(1/\varepsilon))$. For the WPRG case, Pyne and Vadhan [PV21] showed a $\varepsilon$-WPRG with seed length $\widetilde{O}(\log n\sqrt{\log(n/\varepsilon)}+\log(1/\varepsilon))$. This was improved by Chen, Hoza, Lyu, Tal, and Wu [CHL+23] to $O\left(\log n\sqrt{\log(1/\varepsilon)}\sqrt{\log\log(n/\varepsilon)}+\log(1/\varepsilon)\log\log(n/\varepsilon)\right)$.
We give an improved WPRG against permutation ROBPs as the following.
@@
for the class of permutation ROBPs of length $n$ and alphabet $\{0,1\}^{s}$ with a single accept node.
The comparison of our result with the previous results is shown in Table 3.
@@
Note that by replacing $\varepsilon$ with $\varepsilon/w$, one can attain a WPRG for multiple accept nodes. In this way our result gives a WPRG with seed length $O\left(s+\log n\left(\log\log n+\sqrt{\log(w/\varepsilon)}\right)+\log(w/\varepsilon)\right)$. This reveals a substantial difference between our result and [CHL+23]. Notice that our seed length is optimal $O(\log w)$ for $n=2^{O(\sqrt{\log w})},\varepsilon=1/\operatorname{\mathrm{poly}}w$, i.e. it implies a Nisan-Zuckerman style PRG for short-wide permutation ROBPs. The only prior work attaining optimal seed length for such ROBPs are [NZ96, Arm98] but only work for $n=\operatorname{\mathrm{poly}}\log w,\varepsilon=2^{-\log^{0.9}w}$. It is not clear how to use [CHL+23] to achieve optimal seed length even for constant error, since when switching to multiple accept nodes, the only known way is to replace $\varepsilon$ with $\varepsilon/w$.
We also remark that one may want to use the sampler technique of [Hoz21] on the WPRG of [CHL+23] to attain similar parameters to our Theorem 1.9. However, one barrier is that it is unclear if it is possible to attain samplers supporting the sv-approximation used in the analysis of [CHL+23], with desired parameters.
1.4 Derandomization for short-wide regular ROBPs
Our framework also provides a new derandomization for regular ROBPs. There is an extensive body of work on constructing HSGs, WPRGs, and PRGs for regular ROBPs [BRRY14, BHPP22, De11, RSV13, CL24, CHL+23]. For derandomization, dedicated derandomization algorithms achieve better performance. Ahmadinejad, Kelner, Murtagh, Peebles, Sidford, and Vadhan [AKM+20] derandomized regular ROBPs of length $n$, width $w$ and alphabet $\{0,1\}$ within error $\varepsilon$, using space $O(\log(nw)\log\log(nw/\varepsilon))$. Subsequently, Chen, Hoza, Lyu, Tal, and Wu [CHL+23] achieved the same result with a simplified algorithm. Chattopadhyay and Liao [CL24] constructed another alternate algorithm with the same space complexity. These derandomizations can be viewed as Sak-Zhou style derandomizations since they can work for full length with small space. However their time complexity is exponential in their space complexity, hence is not polynomial.
Our new derandomization, stated as the following, can be viewed as a Nisan-Zuckerman style derandomization, since it has optimal logspace complexity for short-wide ROBPs.